# Hyperfocal subalgebras of hyperfocal abelian Frobenius blocks

### v1

@@ 3 Hyperfocal abelian Frobenius blocks
Let us borrow the notation from the paragraphs 2.12 and 3.1. Set $P(A_{\delta}^{D})^{*}=(A_{\delta}^{D})^{*}\cdot\{a_{u,\delta}\mid u\in P\}$. Then $P(A_{\delta}^{D})^{*}$ is a subgroup of $N_{A_{\delta}}(D)$.
By [25, Proposition 2.8], we have
$F_{A}(D)=(P(A_{\delta}^{D})^{*})\cdot N_{\mathbb{A}_{\delta}}(D)/D(A_{\delta}^{D})^{*}$
and $N_{\mathbb{A}_{\delta}}(D)(A_{\delta}^{D})^{*}/D(A_{\delta}^{D})^{*}$ is normal in $F_{A}(D)$. Take an element $\mathfrak{a}_{\delta}$ of $P(A_{\delta}^{D})^{*}\cap N_{\mathbb{A}_{\delta}}(D)$. So there is some $u\in P$ such that ${{}^{\mathfrak{a}_{\delta}}}vj={{}^{u}}vj$ for any $v\in D$. Then $\mathfrak{a}_{\delta}^{-1}u$ is in $(j\mathbb{A}ju)^{D}$. By the proof of Lemma 3.6, there is some $\mathfrak{a}\in(\mathbb{A}^{D})^{*}$ such that $\mathfrak{a}^{-1}j\mathfrak{a}=u^{-1}ju$. Therefore, $j\mathbb{A}ju=j\mathbb{A}u\mathfrak{a}^{-1}j\mathfrak{a}=j\mathbb{A}uj\mathfrak{a}$. Combining these facts together, we have $(j\mathbb{A}uj)(D)\neq 0$. By Lemma 3.6, this forces $u\in DC_{P}(D)$. Then the element $\mathfrak{a}_{\delta}$ has to belong to $D(\mathbb{A}_{\delta}^{D})^{*}$. Hence, $(P(A_{\delta}^{D})^{*}/D(A_{\delta}^{D})^{*})\cap(N_{\mathbb{A}_{\delta}}(D)(A_{\delta}^{D})^{*}/D(A_{\delta}^{D})^{*})=1$.
On the other hand, $F_{A}(D)$ is isomorphism to $N_{G}(D_{\delta})/DC_{G}(D)$ with $PC_{G}(D)/DC_{G}(D)$ being a Sylow $p$-subgroup. It is clear $P(A_{\delta}^{D})^{*}/D(A_{\delta}^{D})^{*}$ is the preimage of $PC_{G}(D)/DC_{G}(D)$ under this isomorphism. In conclusion, $F_{A}(D)$ is a $p$-nilpotent group. So is $N_{G}(D_{\delta})/DC_{G}(D)$. ∎
@@
Now we can define the hyperfocal abelian Frobenius block. Keep the notation above. By the Schur-Zassenhaus Theorem, we can identify $E_{\mathfrak{h}}$ with a subgroup of $\mathrm{Aut}(D)$ and then it can act on $D$ by conjugation. Then the block $b$ is called with Frobenius hyperfocal inertial quotient if $E_{\mathfrak{h}}$ is nontrivial and acts freely on $D-\{1\}$, equivalently, the corresponding semidirect product $D\rtimes E_{\mathfrak{h}}$ is a Frobenius group. Moreover, a block $b$ with Frobenius hyperfocal inertial quotient is called a hyperfocal abelian Frobenius block if a hyperfocal subgroup of the block $b$ is abelian.
@@
There are some examples of hyperfocal abelian Frobenius blocks. The first well-known example is the block with cyclic hyperfocal subgroup. The second example is the block with Klein four hyperfocal subgroup. In this case, by [7, Proposition 2.3], $N_{G}(D_{\delta})/C_{G}(D)$ is isomorphic to a cyclic group of order $3$, or the symmetric group $\mathrm{S}_{3}$ of degree 3, depending on $C_{P}(D)=P$ or not. In both cases, they are both hyperfocal abelian Frobenius blocks.
@@
For any element $x$ of $\mbox{$\mathcal{O}$}G$, we use $\bar{x}$ to denote the image of $x$ in $kG$ under the natural surjective homomorphism. There is an analoguous notation for the subset of $\mbox{$\mathcal{O}$}G$. Denote by $b_{P}$ the block of $C_{G}(P)$ over $\mathcal{O}$ such that the block $\bar{b}_{P}$ of $kC_{G}(P)$ is associated with the local pointed group $P_{\gamma}$. Hence, the pair $(P,\bar{b}_{P})$ is a maximal $b$-Brauer pair. It is well-known that for any subgroup $R$ of $P$ there is a unique block $b_{R}$ of $C_{G}(R)$ over $\mathcal{O}$ with $(R,\bar{b}_{R})\leq(P,\bar{b}_{P})$ and the pair $(R,\bar{b}_{R})$ is called a $b$-Brauer pair. Denote by $\mathcal{B}_{G}(b)_{\leq(P,\bar{b}_{P})}$ the Brauer category of the block $b$ with objects consisting of all $b$-Brauer pairs contained in $(P,\bar{b}_{P})$. Let $H$ be a subgroup of $G$. $H$ is said to control fusion in $\mathcal{B}_{G}(b)_{\leq(P,\bar{b}_{P})}$, or control fusion of the block $b$ for simplicity, if the morphisms of the Brauer category $\mathcal{B}_{G}(b)_{\leq(P,\bar{b}_{P})}$ can be induced by elements of $H$ (see [29, §49]).
@@
Let us recall the definitions of some special local pointed groups. A self-centralizing pointed group $Q_{\epsilon}$ on $\mbox{$\mathcal{O}$}G$ is a local pointed group on $\mbox{$\mathcal{O}$}G$ such that $Z(Q)$ is a defect group of the block $b(\epsilon)$ of $C_{G}(Q)$ associated with $Q_{\epsilon}$. Furthermore, a self-centralizing pointed group $Q_{\epsilon}$ is called an essential pointed group if the quotient $N_{G}(Q_{\epsilon})/QC_{G}(Q)$ contains a proper subgroup $M$ such that $p$ divides $|M|$ but does not divide $|M\cap M^{x}|$ for any $x\in(N_{G}(Q_{\epsilon})/QC_{G}(Q))-M$. In particular, $N_{G}(Q_{\epsilon})/QC_{G}(Q)$ has no nontrivial normal $p$-subgroup. For an essential pointed group $Q_{\epsilon}$, it is clear that $N_{G}(Q_{\epsilon})=N_{G}(Q,\bar{b}(\epsilon))$ and we also call the corresponding $b$-Brauer pair $N_{G}(Q,\bar{b}(\epsilon))$ an essential $b$-Brauer pair.
@@
Let $Q$ be a nontrivial subgroup of $D$. Suppose that $C_{P}(Q)$ is a defect group of the block $b_{Q}$ of $C_{G}(Q)$. We denote $C_{P}(Q)$ by $Q_{c}$. We first prove this proposition for this special case. For any subgroup $R$ of $C_{P}(Q)$, $(R,\bar{b}_{QR})$ is a $b_{Q}$-Brauer pair. In particular $(Q_{c},\bar{b}_{Q_{c}})$ is a maximal $b_{Q}$-Brauer pair. In order to show that the block $b_{Q}$ is nilpotent, it suffices to prove that the hyperfocal subgroup of the block $b_{Q}$ is trivial. By [25, 1.7.1], it reduces to calculate that $N_{C_{G}(Q)}(R,\bar{b}_{QR})/C_{G}(QR)$ is a $p$-group when $(R,\bar{b}_{QR})$ is an essential $b_{Q}$-Brauer pair and $(R,\bar{b}_{QR})=(Q_{c},\bar{b}_{Q_{c}})$.
@@
Obviously, $N_{G}(D,\bar{b}_{D})/C_{G}(D)$ is isomorphic to $N_{G}(D_{\delta})/C_{G}(D)$. Since $E_{\mathfrak{h}}$ acts freely on $D-\{1\}$, we have $N_{G}(D,\bar{b}_{D})\cap C_{G}(Q)=Q_{c}C_{G}(D)$.
@@
In general, it is well-known that there is some element $x$ of $G$ such that $(Q,\bar{b}_{Q})\leq(P,\bar{b}_{P})^{x}$ and $C_{P^{x}}(Q)$ is a defect group of the block $b_{Q}$. Since the hyperfocal subgroup $D$ is abelian, we can assume that $x$ is in $N_{G}(D,\bar{b}_{D})$ by [33, Theorem 2]. Therefore, $C_{P}({{}^{x}}Q)$ is a defect group of the block ${{}^{x}}b_{Q}$ of $C_{G}({{}^{x}}Q)$. Since $x$ belongs to $N_{G}(D,\bar{b}_{D})$, ${{}^{x}}Q$ is still a nontrivial subgroup of $D$ and $({{}^{x}}Q,{{}^{x}}\bar{b}_{Q})$ is contained in $(D,\bar{b}_{D})$. By the uniquness of $\bar{b}_{{{}^{x}}Q}$, we get ${{}^{x}}b_{Q}=b_{{{}^{x}}Q}$. On the other hand, by the argument for the special case, the block $b_{{{}^{x}}Q}$ of $C_{G}({{}^{x}}Q)={{}^{x}}C_{G}(Q)$ is nilpotent. Obviously, the conjugation action of $x$ induces an isomorphism between the Brauer categories of the block $b_{Q}$ and $b_{{{}^{x}}Q}$. Hence, the block $b_{Q}$ of $C_{G}(Q)$ is also nilpotent. ∎
@@ 4 The bimodule structure and the Brauer quotients of $\mathbb{A}_{\delta}$
Keep the notation in the last section. We will first describe the $\mbox{$\mathcal{O}$}(D\times D)$-module structure of $\mathbb{A}_{\delta}$ without assuming that the block $b$ is a hyperfocal abelian Frobenius block. This is analogous to the $\mbox{$\mathcal{O}$}(P\times P)$-module structure of the source algebra $A$ (see [29, Theorem (44.3)]). Let us collect some basic results about the $\mbox{$\mathcal{O}$}(D\times D)$-module structure of $\mathbb{A}_{\delta}$.
@@
The direct summand of $\mathbb{A}_{\delta}$ as an $\mbox{$\mathcal{O}$}(D\times D)$-module has the form $\mbox{$\mathcal{O}$}DgD$ for some $g\in G$, which is isomorphic to $\mathrm{Ind}_{\Delta_{g}(D)}^{D\times D}(\mbox{$\mathcal{O}$})$. Here, $\Delta_{g}(D)=\{(u,g^{-1}ug)\in D\times D\mid u\in D\cap{{}^{g}}D\}$ is a subgroup of $D\times D$. Moreover, setting $Q=D\cap{{}^{g}}D$, then $(Q,\bar{b}_{Q})^{g}\subseteq(D,\bar{b}_{D})$.
@@
Obviously, as $\mbox{$\mathcal{O}$}(D\times D)$-modules, $\mathbb{A}$ is a direct summand of $A$ and $\mathbb{A}_{\delta}$ is a direct summand of $\mathbb{A}$. Therefore, $\mathbb{A}_{\delta}$ is a direct summand of $A$ as $\mbox{$\mathcal{O}$}(D\times D)$-modules. The statements in this lemma can follow from [17, Theorem 8.7.1]. ∎
@@
The existence of a $(D\times D)$-invariant $\mathcal{O}$-basis $\mathcal{B}$ follows from Lemma 4.1. In order to show that the $\mathcal{O}$-basis $\mathcal{B}$ can be chosen to contain $j$, it suffices to demonstrate that this $\mathcal{B}$ contains an invertible element of $\mathbb{A}_{\delta}^{D}$. So we can assume that $\mbox{$\mathcal{O}$}=k$. Set $j=\sum\limits_{x\in\mathcal{B}}\lambda_{x}x$ with $\lambda_{x}\in k$. Then $\mathrm{Br}_{D}(j)=\sum_{x\in\mathcal{B}^{D}}\lambda_{x}\mathrm{Br}_{D}(x)$ is the unit element of $\mathbb{A}_{\delta}(D)$ which is isomorphic to the local algebra $kZ(D)$ by Lemma 3.6. In particular, there is at least one element $x_{0}\in\mathcal{B}^{D}$ with $\lambda_{x_{0}}\neq 0$ and $\mathrm{Br}_{D}(x_{0})$ being invertible in $\mathbb{A}_{\delta}(D)$. Therefore, $x_{0}$ is invertible in $\mathbb{A}_{\delta}^{D}$ by [29, Theorem (3.2)]. ∎
@@
Denote by $\Delta(D)$ the diagonal group of $D\times D$. Then by the Mackey decomposition formula, the direct summand of $\mathrm{Res}^{D\times D}_{\Delta(D)}(\mbox{$\mathcal{O}$}DgD)$ has the form $\mathrm{Ind}_{\Delta(D)\cap{{}^{(u,v)}}\Delta_{g}(D)}^{\Delta(D)}(\mbox{$\mathcal{O}$})$. It is easy to check that $\Delta(D)\cap{{}^{(u,v)}}\Delta_{g}(D)=\Delta(D)$ if and only if $g\in DC_{G}(D)$ and when $g\in DC_{G}(D)$, $\mbox{$\mathcal{O}$}DgD=\mbox{$\mathcal{O}$}Dg$ is isomorphic to $\mbox{$\mathcal{O}$}D$ as $\mbox{$\mathcal{O}$}(D\times D)$-modules. Hence, $(\mbox{$\mathcal{O}$}DgD)(D)$ is isomorphic to $kZ(D)$ when $g\in DC_{G}(D)$. By Lemma 3.6, we have $\mathbb{A}_{\delta}(D)\cong kZ(D)$. Therefore, there is only one direct summand of $\mathbb{A}_{\delta}$ isomorphic to $\mbox{$\mathcal{O}$}D$. ∎
@@
Recall that $N_{G}(D_{\delta})/DC_{G}(D)$ is a $p$-nilpotent group with normal $p$-complement $E_{\mathfrak{h}}$. By the proof of Proposition 3.7, the isomorphism from $N_{G}(D_{\delta})/DC_{G}(D)$ to $F_{A}(D)$ maps $E_{\mathfrak{h}}$ onto $N_{\mathbb{A}_{\delta}}(D)(A_{\delta}^{D})^{*}/D(A_{\delta}^{D})^{*}$, which is isomorphic to $N_{\mathbb{A}_{\delta}}(D)/D(\mathbb{A}_{\delta}^{D})^{*}$. For any $g\in N_{G}(D_{\delta})$ with $gDC_{G}(D)\in E_{\mathfrak{h}}$, let $\mathfrak{a}_{g}\in\mathbb{A}_{\delta}^{*}$ such that $\mathfrak{a}_{g}(uj)\mathfrak{a}_{g}^{-1}=(gug^{-1})j$ for any $u\in D$.
@@
It is easy to check that mapping $u\mathfrak{a}_{g}$ to $ug$ gives an $\mbox{$\mathcal{O}$}(D\times D)$-module isomorphism from $\mbox{$\mathcal{O}$}D\mathfrak{a}_{g}$ to $\mbox{$\mathcal{O}$}Dg$. By Lemma 4.2, we can write $\mathbb{A}_{\delta}=\mbox{$\mathcal{O}$}Dj\oplus Y$ as $\mbox{$\mathcal{O}$}(D\times D)$-modules for some $\mbox{$\mathcal{O}$}(D\times D)$-submodule $Y$. Then we have $\mathbb{A}_{\delta}=\mathbb{A}_{\delta}\mathfrak{a}_{g}=\mbox{$\mathcal{O}$}D\mathfrak{a}_{g}\oplus Y\mathfrak{a}_{g}$. Hence, $\mbox{$\mathcal{O}$}D\mathfrak{a}_{g}$ is a direct summand of $\mathbb{A}_{\delta}$. By Lemma 4.3, we can get the multiplicity of $\mbox{$\mathcal{O}$}D\mathfrak{a}_{g}$ is $1$. ∎

### v2

@@ 3 Hyperfocal abelian Frobenius blocks
Set $P(A_{\delta}^{D})^{\times}=(A_{\delta}^{D})^{\times}\cdot\{a_{u,\delta}\mid u\in P\}$. It is easy to check that $P(A_{\delta}^{D})^{\times}/D(A_{\delta}^{D})^{\times}$ is just the image of $PC_{G}(D)/DC_{G}(D)$, which is a Sylow $p$-subgroup of $N_{G}(D_{\delta})/DC_{G}(D)$, under the isomorphism $\mathfrak{f}_{G,D}$. Hence, $P(A_{\delta}^{D})^{\times}/D(A_{\delta}^{D})^{\times}$ is a Sylow $p$-subgroup of $F_{A}(D)$.
On the other hand, by [25, Proposition 2.8], we have
$F_{A}(D)=((P(A_{\delta}^{D})^{\times})\cdot N_{\mathbb{A}_{\delta}}(D))/D(A_{\delta}^{D})^{\times}$
and $N_{\mathbb{A}_{\delta}}(D)(A_{\delta}^{D})^{\times}/D(A_{\delta}^{D})^{\times}$ is normal in $F_{A}(D)$. Take an element $\mathfrak{a}_{\delta}$ of $P(A_{\delta}^{D})^{\times}\cap N_{\mathbb{A}_{\delta}}(D)$. So there is some $u\in P$ such that ${{}^{\mathfrak{a}_{\delta}}}(vj)=({{}^{u}}v)j$ for any $v\in D$. Then $\mathfrak{a}_{\delta}^{-1}u$ is in $(j\mathbb{A}ju)^{D}$ and ${\rm Br}_{D}(\mathfrak{a}_{\delta}^{-1}u)\neq 0$ since $\mathfrak{a}_{\delta}^{-1}u\cdot u^{-1}\mathfrak{a}_{\delta}=j$. By the paragraph 3.1, there is some $a_{u}\in(\mathbb{A}^{D})^{\times}$ such that $a_{u}^{-1}ja_{u}=u^{-1}ju$. Therefore, $j\mathbb{A}ju=j\mathbb{A}ua_{u}^{-1}ja_{u}=j\mathbb{A}uja_{u}$. Combining these facts together, we have $(j\mathbb{A}uj)(D)\neq 0$. By Lemma 3.6, this forces $u\in DC_{P}(D)$. Then the element $\mathfrak{a}_{\delta}$ has to belong to $D(\mathbb{A}_{\delta}^{D})^{\times}$. Hence, $(P(A_{\delta}^{D})^{\times}/D(A_{\delta}^{D})^{\times})\cap(N_{\mathbb{A}_{\delta}}(D)(A_{\delta}^{D})^{\times}/D(A_{\delta}^{D})^{\times})=1$. In particular, $F_{A}(D)$ is a $p$-nilpotent group. So is $N_{G}(D_{\delta})/DC_{G}(D)$. The last statement can be easily obtained. ∎
@@
Now we can define the hyperfocal abelian Frobenius block. Keep the notation above. By the Schur-Zassenhaus Theorem, we can identify $E_{\mathfrak{h}}$ with a subgroup of $\mathrm{Aut}(D)$ and then it can act on $D$ by the conjugation action. Then the block $b$ is called with Frobenius hyperfocal inertial quotient if $D$ and $E_{\mathfrak{h}}$ are both nontrivial and $E_{\mathfrak{h}}$ acts freely on $D-\{1\}$, equivalently, $D$ is nontrivial and the corresponding semidirect product $D\rtimes E_{\mathfrak{h}}$ is a Frobenius group. Moreover, a block $b$ with Frobenius hyperfocal inertial quotient is called a hyperfocal abelian Frobenius block if a hyperfocal subgroup of the block $b$ is abelian.
@@
There are some examples of hyperfocal abelian Frobenius blocks. The first well-known example is the block with a cyclic hyperfocal subgroup. The second example is the block with a Klein four hyperfocal subgroup. In this case, by [7, Proposition 2.3], $N_{G}(D_{\delta})/C_{G}(D)$ is isomorphic to a cyclic group of order $3$, or the symmetric group $\mathrm{S}_{3}$ of degree 3, depending on $C_{P}(D)=P$ or not. In both cases, they are both hyperfocal abelian Frobenius blocks.
@@
For any element $x$ of $\mbox{$\mathcal{O}$}G$, we use $\bar{x}$ to denote the image of $x$ in $kG$ under the natural surjective homomorphism. There is an analogous notation for the subset of $\mbox{$\mathcal{O}$}G$. First we fix some notation about $b$-Brauer pairs. We refer to [29, §40] for details. Denote by $b_{P}$ the block of $C_{G}(P)$ over $\mathcal{O}$ such that the block $\bar{b}_{P}$ of $kC_{G}(P)$ is associated with the local pointed group $P_{\gamma}$. Then the pair $(P,\bar{b}_{P})$ is a maximal $b$-Brauer pair. For any subgroup $R$ of $P$, let $b_{R}$ be the unique block of $C_{G}(R)$ over $\mathcal{O}$ such that $(R,\bar{b}_{R})\leq(P,\bar{b}_{P})$. Denote by $\mathcal{B}_{G}(b)_{\leq(P,\bar{b}_{P})}$ the Brauer category of the block $b$ whose objects consist of all $b$-Brauer pairs contained in $(P,\bar{b}_{P})$ and morphisms from $(Q,\bar{b}_{Q})$ to $(R,\bar{b}_{R})$ are group homomorphisms from $Q$ to $R$ induced by the conjugation action of an element $g$ in $G$ such that ${{}^{g}}(Q,\bar{b}_{Q})\leq(R,\bar{b}_{R})$. Let $H$ be a subgroup of $G$. $H$ is said to control fusion in $\mathcal{B}_{G}(b)_{\leq(P,\bar{b}_{P})}$, or control fusion of the block $b$ for simplicity, if the morphisms of the Brauer category $\mathcal{B}_{G}(b)_{\leq(P,\bar{b}_{P})}$ can be induced by elements in $H$ (see [29, §49]).
@@
Let us recall the definitions of some special local pointed groups (see [25, 1.6]). A self-centralizing pointed group $Q_{\epsilon}$ on $\mbox{$\mathcal{O}$}G$ is a local pointed group on $\mbox{$\mathcal{O}$}G$ such that $Z(Q)$ is a defect group of the block $\bar{b}(\epsilon)$ of $C_{G}(Q)$ associated with $Q_{\epsilon}$. Furthermore, a self-centralizing pointed group $Q_{\epsilon}$ is called an essential pointed group if the quotient $N_{G}(Q_{\epsilon})/QC_{G}(Q)$ contains a proper subgroup $M$ such that $p$ divides $|M|$ but does not divide $|M\cap M^{x}|$ for any $x\in N_{G}(Q_{\epsilon})/QC_{G}(Q)-M$. In particular, $N_{G}(Q_{\epsilon})/QC_{G}(Q)$ has no nontrivial normal $p$-subgroup. For an essential pointed group $Q_{\epsilon}$, it is clear that $N_{G}(Q_{\epsilon})=N_{G}(Q,\bar{b}(\epsilon))$ and we also call the corresponding $b$-Brauer pair $(Q,\bar{b}(\epsilon))$ an essential $b$-Brauer pair.
@@
Let $Q$ be a nontrivial subgroup of $D$. Suppose that $C_{P}(Q)$ is a defect group of the block $b_{Q}$ of $C_{G}(Q)$. We denote $C_{P}(Q)$ by $Q_{c}$. We first prove this proposition for this special case. For any subgroup $R$ of $C_{P}(Q)$, $(R,\bar{b}_{QR})$ is a $b_{Q}$-Brauer pair. In particular $(Q_{c},\bar{b}_{Q_{c}})$ is a maximal $b_{Q}$-Brauer pair. In order to show that the block $b_{Q}$ is nilpotent, it suffices to prove that the hyperfocal subgroup of the block $b_{Q}$ is trivial. By [25, 1.7.1], it reduces to show that $N_{C_{G}(Q)}(R,\bar{b}_{QR})/C_{G}(QR)$ is a $p$-group when $(R,\bar{b}_{QR})$ is either an essential $b_{Q}$-Brauer pair or $(R,\bar{b}_{QR})=(Q_{c},\bar{b}_{Q_{c}})$.
@@
Obviously, $N_{G}(D,\bar{b}_{D})/C_{G}(D)$ is isomorphic to $N_{G}(D_{\delta})/C_{G}(D)$. Since $E_{\mathfrak{h}}$ acts freely on $D-\{1\}$ and $D$ is abelian, we have $N_{G}(D,\bar{b}_{D})\cap C_{G}(Q)=Q_{c}C_{G}(D)$.
@@
In general, it is well-known that there is some element $x$ of $G$ such that $(Q,\bar{b}_{Q})\leq(P,\bar{b}_{P})^{x}$ and $C_{P^{x}}(Q)$ is a defect group of the block $b_{Q}$. Since the hyperfocal subgroup $D$ is abelian, we can assume that $x$ is in $N_{G}(D,\bar{b}_{D})$ by [33, Theorem 2]. Therefore, $C_{P}({{}^{x}}Q)$ is a defect group of the block ${{}^{x}}b_{Q}$ of $C_{G}({{}^{x}}Q)$. Since $x$ belongs to $N_{G}(D,\bar{b}_{D})$, ${{}^{x}}Q$ is still a nontrivial subgroup of $D$ and $({{}^{x}}Q,{{}^{x}}\bar{b}_{Q})$ is contained in $(D,\bar{b}_{D})$. By the uniquness of $\bar{b}_{{{}^{x}}Q}$, we get ${{}^{x}}b_{Q}=b_{{{}^{x}}Q}$. On the other hand, by the argument above for the special case, the block $b_{{{}^{x}}Q}$ of $C_{G}({{}^{x}}Q)={{}^{x}}C_{G}(Q)$ is nilpotent. Obviously, the conjugation action of $x$ induces an isomorphism between the Brauer categories of the block $b_{Q}$ and $b_{{{}^{x}}Q}$. Hence, the block $b_{Q}$ of $C_{G}(Q)$ is also nilpotent. ∎
@@ 4 The bimodule structure and the Brauer quotients of $\mathbb{A}_{\delta}$
Keep the notation in the last section. In this sectoin, we will first describe the $\mbox{$\mathcal{O}$}(D\times D)$-module structure of $\mathbb{A}_{\delta}$ without assuming that the block $b$ is a hyperfocal abelian Frobenius block. This is analogous to the $\mbox{$\mathcal{O}$}(P\times P)$-module structure of the source algebra $A$ (see [29, Theorem (44.3)]). Let us collect some basic results about the $\mbox{$\mathcal{O}$}(D\times D)$-module structure of $\mathbb{A}_{\delta}$.
@@
Every indecomposable direct summand of the $\mbox{$\mathcal{O}$}(D\times D)$-module $\mathbb{A}_{\delta}$ has the form $\mbox{$\mathcal{O}$}DgD$ for some $g\in G$, which is isomorphic to $\mathrm{Ind}_{\Delta_{g}(D)}^{D\times D}(\mbox{$\mathcal{O}$})$. Here, $\Delta_{g}(D)=\{(u,g^{-1}ug)\in D\times D\mid u\in D\cap{{}^{g}}D\}$ is a subgroup of $D\times D$. Moreover, setting $Q=D\cap{{}^{g}}D$, then $(Q,\bar{b}_{Q})^{g}\subseteq(D,\bar{b}_{D})$.
@@
Obviously, as $\mbox{$\mathcal{O}$}(D\times D)$-modules, $\mathbb{A}$ is a direct summand of $A$ and $\mathbb{A}_{\delta}$ is a direct summand of $\mathbb{A}$. Therefore, $\mathbb{A}_{\delta}$ is isomorphic to a direct summand of $A$ as $\mbox{$\mathcal{O}$}(D\times D)$-modules. The statements in this lemma can follow from [16, Theorem 8.7.1]. ∎
@@
The existence of a $(D\times D)$-invariant $\mathcal{O}$-basis $\mathcal{B}$ follows from Lemma 4.1. In order to show that the $\mathcal{O}$-basis $\mathcal{B}$ can be chosen to contain $j$, it suffices to demonstrate that this $\mathcal{B}$ contains an invertible element in $\mathbb{A}_{\delta}^{D}$. So we can assume that $\mbox{$\mathcal{O}$}=k$. Set $j=\sum\limits_{x\in\mathcal{B}}\lambda_{x}x$ with $\lambda_{x}\in k$. Then $\mathrm{Br}_{D}(j)=\sum\limits_{x\in\mathcal{B}^{D}}\lambda_{x}\mathrm{Br}_{D}(x)$ is the unit element of $\mathbb{A}_{\delta}(D)$ which is isomorphic to the local algebra $kZ(D)$ by Lemma 3.6. In particular, there is at least one element $x_{0}\in\mathcal{B}^{D}$ with $\lambda_{x_{0}}\neq 0$ and $\mathrm{Br}_{D}(x_{0})$ being invertible in $\mathbb{A}_{\delta}(D)$. Therefore, $x_{0}$ is invertible in $\mathbb{A}_{\delta}^{D}$ by [29, Theorem (3.2)]. ∎
@@
Denote by $\Delta(D)$ the diagonal group of $D\times D$. Then by Mackey’s formula, the direct summand of $\mathrm{Res}^{D\times D}_{\Delta(D)}(\mbox{$\mathcal{O}$}DgD)$ has the form $\mathrm{Ind}_{\Delta(D)\cap{{}^{(u,v)}}\Delta_{g}(D)}^{\Delta(D)}(\mbox{$\mathcal{O}$})$ with $(u,v)\in D\times D$. It is easy to check that $\Delta(D)\cap{{}^{(u,v)}}\Delta_{g}(D)=\Delta(D)$ for some $(u,v)\in D\times D$ if and only if $g\in DC_{G}(D)$ and when $g\in DC_{G}(D)$, $\mbox{$\mathcal{O}$}DgD=\mbox{$\mathcal{O}$}Dg$ is isomorphic to $\mbox{$\mathcal{O}$}D$ as $\mbox{$\mathcal{O}$}(D\times D)$-modules. Hence, $(\mbox{$\mathcal{O}$}DgD)(D)$ is isomorphic to $kZ(D)$ when $g$ belongs to $DC_{G}(D)$. Then by Lemma 3.6, there is only one indecomposable direct summand of the $\mbox{$\mathcal{O}$}(D\times D)$-module $\mathbb{A}_{\delta}$ isomorphic to $\mbox{$\mathcal{O}$}D$. ∎
@@
Recall that $N_{G}(D_{\delta})/DC_{G}(D)$ is a $p$-nilpotent group with normal $p$-complement $E_{\mathfrak{h}}$. By Proposition 3.7, the isomorphism $\mathfrak{f}_{G,D}$ from $N_{G}(D_{\delta})/DC_{G}(D)$ to $F_{A}(D)$ maps $E_{\mathfrak{h}}$ onto $N_{\mathbb{A}_{\delta}}(D)(A_{\delta}^{D})^{\times}/D(A_{\delta}^{D})^{\times}$, which is isomorphic to $N_{\mathbb{A}_{\delta}}(D)/D(\mathbb{A}_{\delta}^{D})^{\times}$. For any $g\in N_{G}(D_{\delta})$ with $gDC_{G}(D)\in E_{\mathfrak{h}}$, we denote by $\mathfrak{a}_{g}$ an element in $\mathbb{A}_{\delta}^{\times}$ such that $\mathfrak{a}_{g}(uj)\mathfrak{a}_{g}^{-1}=(gug^{-1})j$ for any $u\in D$.
@@
It is easy to check that mapping $u\mathfrak{a}_{g}$ to $ug$ gives an $\mbox{$\mathcal{O}$}(D\times D)$-module isomorphism from $\mbox{$\mathcal{O}$}D\mathfrak{a}_{g}$ to $\mbox{$\mathcal{O}$}Dg$. By Lemma 4.2, we can write $\mathbb{A}_{\delta}=\mbox{$\mathcal{O}$}Dj\oplus Y$ as $\mbox{$\mathcal{O}$}(D\times D)$-modules for some $\mbox{$\mathcal{O}$}(D\times D)$-submodule $Y$. Then we have $\mathbb{A}_{\delta}=\mathbb{A}_{\delta}\mathfrak{a}_{g}=\mbox{$\mathcal{O}$}D\mathfrak{a}_{g}\oplus Y\mathfrak{a}_{g}$ which is still a decomposition of $\mathbb{A}_{\delta}$ as $\mbox{$\mathcal{O}$}(D\times D)$-modules. Hence, $\mbox{$\mathcal{O}$}D\mathfrak{a}_{g}$ is a direct summand of $\mathbb{A}_{\delta}$. By Lemma 4.3, we can get that the multiplicity of $\mbox{$\mathcal{O}$}D\mathfrak{a}_{g}$ is $1$. ∎